OpenStudy (anonymous):

DeJuana bought a lottery ticket and used her birthday to pick the numbers: 6 2 8 0. What is the probability that she has the winning ticket if the possible numbers are 0 through 9? A.1/6561 B.1/40 C.1/36 D.1/10000

Directrix (directrix):

There are 10 digits from 0 through 9. The probability of a six is (1/10) as is the probability of a "2," "8," and "0." Assumption: for a winning ticket for Dejuana, the digits must fall in the order 6-2-8-0. P(6280) = (1/10)* (1/10) * (1/10) * (1/10) = 1/10000

OpenStudy (anonymous):

thanks.... thats helpful